OpenStudy (ilovepuppies):

Check my answer P(10)+P(others) = 16 P(16)-P(10) = 6 6 - 10 = 1:4

OpenStudy (aaronandyson):

Are you taking their ratios?

OpenStudy (ilovepuppies):

I don't know what the means but. P(10)+P(1+5) = 16 So 16 - 10 = 6 That leaves me with 6-10= 4 which is 1:4

OpenStudy (ilovepuppies):

@mathmate can you check my answer?

OpenStudy (ilovepuppies):

@mathmate

OpenStudy (mathmate):

Probabilities never exceed 1! For the ticket question, you don't have to find probabilities to find odds. For \(\color{red}{example}\), the odds \(for\) winning a green tickets = # of green tickets : # of non-green tickets = 100 (green) : 25 (yellow) +5(blue) =100 : 30 = 10 : 3 (after simplifications)

OpenStudy (mathmate):

@ILovePuppies It's not a good idea to start a new post for the same question without a reference. Other people will not be able to follow your thoughts. Even with a reference, it is hard for others to go back and forth between two posts! Thank you for understanding.

OpenStudy (ilovepuppies):

Ok thank you